Flint Corporation issued 2,000 $1,000 bonds at 102. Each bond was issued with one detachable stock warrant. After issuance, the bonds were selling in the market at 98, and the warrants had a market price of $39.

Use the proportional method to record the issuance of the bonds and warrants. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Round intermediate calculations to 5 decimal places, e.g. 1.24687 and final answers to 0 decimal places, e.g. 5,125.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount
Answer #1

Journal Entry

No Account Title and Explanation Debit Credit
1 Cash $2,040,000
Discount on Bonds Payable $38,077
Bonds Payable $2000,000
Paid in Capital : Stock Warrants $78,077

Cash = 2,000 × 1,000 × 102 % = $2,040,000

Discount on Bonds Payable = $2,000,000 - $1,961,923 = $38,077

$2,040,000 ×980 / (980 +39) = $1,961,923